# Interpreting Graph Inference with Skyline Explanations

### v3

@@ I Introduction
Graph models such as graph neural networks ($\mathsf{GNNs}$) have been trained and routinely queried to perform network analysis in, e.g., biochemistry, social and financial networks [1, 2, 3, 4], among other graph analytical tasks. Given a graph $G$ and a set of test nodes $V_{T}$ in $G$, a $\mathsf{GNN}$ $\mathcal{M}$} can be considered as a function that converts a feature representation of $G$ in the form of $(X,A)$, where $X$ (resp. $A$) refers to the node feature matrix (resp. normalized adjacency matrix) of $G$, to an output representation $Z$ (“output embedding matrix”). The output $Z$ can be post-processed to task-specific output such as labels for classification. An “inference query” invokes the above process to query $\mathcal{M}$}, to perform the inference analysis of $G$.
Despite the promising performance of $\mathsf{GNNs}$, it remains desirable yet nontrivial to interpret their outputs to help users understand $\mathsf{GNN}$-based decision making [5]. Several methods (“explainers”) have been proposed  [6, 7, 8, 9, 10]. Typically, a $\mathsf{GNN}$ explainer extracts a subgraph $G_{\zeta}$ of $G$ that can best clarify the output of an inference query posed on $\mathcal{M}$} over $G$. This is often addressed by discovering a subgraph $G_{\zeta}$ subject to a pre-defined metric, which quantifies the explainability of $G_{\zeta}$ for the output (as summarized in Table I).
For example, a subgraph $G_{\zeta}$ of a graph $G$ is a “factual” explanation for $\mathcal{M}$} over $G$, if it preserves the output of $\mathcal{M}$} (hence is “faithful” to the output of $\mathcal{M}$}[6, 10, 11, 8]. $G_{\zeta}$ is a “counterfactual” explanation, if removing the edges of $G_{\zeta}$ from $G$ leads to a change of the output of $\mathcal{M}$} on the remaining graph (denoted as $G\setminus G_{\zeta}$[7, 11, 8]. Other metrics include $\mathsf{fidelity}$$^{-}$ [7, 11, 8] (resp. $\mathsf{fidelity}$$^{+}$ [10, 12]), which quantifies the explainability of $G_{\zeta}$ in terms of the closeness between the task-specific output of $\mathcal{M}$}, such as the probability of label assignments, over $G_{\zeta}$ (resp. $G\setminus G_{\zeta}$) and their original counterpart $G$, and “conciseness (sparsity)” that favors small explanatory subgraphs.
@@
Figure 1: A Bitcoin transaction network with a target IP address (test node $v_{t}$) that has a label “illicit” to be explained [13].
@@
$\mathsf{fac}$ factual ${\mathcal{M}}(v,G)$ = ${\mathcal{M}}(v,G_{\zeta})?$ $\mathsf{true}$$\mathsf{false}$ } a Boolean function  [6, 10, 11, 8]
$\mathsf{cfac}$ counterfactual ${\mathcal{M}}(v,G)\neq{\mathcal{M}}(v,G\setminus G_{\zeta})?$ $\mathsf{true}$$\mathsf{false}$ } a Boolean function  [7, 11, 8]
$\mathsf{fdl}$$^{+}$ fidelity+ $Pr({\mathcal{M}}(v,G))-Pr({\mathcal{M}}(v,G\setminus G_{\zeta}))$ [1,1][-1,1] the larger, the better  [14, 12]
$\mathsf{fdl}$$^{-}$ fidelity- $Pr({\mathcal{M}}(v,G))-Pr({\mathcal{M}}(v,G_{\zeta}))$ [1,1][-1,1] the smaller, the better  [14, 10]
$\mathsf{conc}$ conciseness $\frac{1}{N}\sum_{i=1}^{N}(1-\frac{|E(G_{\zeta})|}{|E(G)|})$ $[0,1]$ the smaller, the better  [10, 12]
$\mathsf{shapley}$ Shapley value $\phi(G_{\zeta})=\sum_{S\subseteq P\setminus\{G_{\zeta}\}}\frac{|S|!(|P|-|S|-1)!}{|P|!}m(S,G_{\zeta})$ [1,1][-1,1] total contribution of nodes in $G_{\zeta}$  [12]
@@
A law enforcement agency has posed an inference query over $\mathcal{M}$} to get the output, and wants to further understand “why” the $\mathsf{GNN}$ asserts $v_{t}$ as an illicit account address. They may further ask clarifications by asking “which fraction of the graph $G$ is responsible for the $\mathsf{GNN}$’s decision on assigning the label “illicit” to the account $v_{t}$?”, and ground this output by referring to known real-world money laundering scenarios [15, 16]). For example, “Spindle” [15] suggests that perpetrators generate multiple shadow addresses to transfer small amounts of assets along lengthy paths to a specific destination; and Peel Chain [16] launders large amounts of cryptocurrency through sequences of small transactions, where minor portions are ‘peeled’ from the original address and sent for conversion into fiat currency to minimize the risk.
An explanatory subgraph $G_{\zeta}^{1}$ is a factual explanatory subgraph generated by the explainer in [6]; and $G_{\zeta}^{2}$ is a counterfactual one generated by [7]. Comparing $G_{\zeta}^{1}$ with $G_{\zeta}^{2}$, we observe the following. (1) $G_{\zeta}^{1}$ is a small and concise explanation that only involves $v_{t}$ and most of its neighbors, which includes neighborhood features that preserve the output of a $\mathsf{GNN}$’s output; nevertheless, it misses important nodes that have a higher influence on $\mathsf{GNN}$’s output, that are not necessarily in $v_{t}$’s direct neighborhood. (2) Such nodes can be captured by a counterfactual explanatory subgraph, as depicted in $G_{\zeta}^{2}$. However, a larger fraction of $G$ is included to ensure that the “removal” of edges incurs enough impact to change the output of the $\mathsf{GNN}$, hence sacrificing “conciseness”, conflicting with users’ need on quick generation of small evidence which is “faithful” to the original $\mathsf{GNN}$ output [11].
@@
Can we generate explanations that simultaneously address multiple explainability metrics? A quick solution is to compute subgraphs that optimize a weighted linear combination of all metrics [17, 18]. However, this may lead to a marginally optimal answer over all measurements, overlooking other high-quality and diverse solutions, hence an overkill.
@@
Example 2: Consider another two explanatory subgraphs: $G_{\zeta}^{3}$ is the explanation generated by [10], an explainer that optimizes both conciseness and factual measures [10]; $G_{\zeta}^{4}$ is from an explainer that linearly combines factual and counterfactual measures into a single, bi-criteria objective function to be optimized [8]. Such methods enforce to optimize potentially “conflicting” measures, or highly correlated ones, either may result in lower-quality solutions that are sensitive to data bias. For example, as fidelity- and conciseness both encourage smaller and faithful explanations, explanatory subgraphs obtained by  [10] such as $G_{\zeta}^{3}$ turns out to be relatively much smaller and less informative, hardly providing sufficient real-world evidence that can be used to interpret money laundering behaviors. On the other hand, explanations from [8] such as $G_{\zeta}^{4}$ may easily be a one-sided, factual or counterfactual evidence. We found that such explanations capture only one type of money laundering scenario at best in most cases. $\Box$
@@
Skyline queries [19, 20, 21, 22, 23] compute Pareto sets (“skylines”) that dominate the rest data points across a set of quality measures. Skylines generally offer better and more comprehensive solutions against the aforementioned alternatives [18, 24]. We advocate approaching $\mathsf{GNN}$ explanation by generating a set of subgraphs that are Poreto sets over multiple user-defined explanatory measures, referred to as “skyline explanations”.
@@
Example 3: Consider a “skyline query” that seeks subgraphs addressing multiple explainability criteria, which returns a set of subgraphs $G_{\zeta}^{5}$, $G_{\zeta}^{6}$, and $G_{\zeta}^{7}$. These explanatory subgraphs are selected as a Pareto-optimal set across three explanatory measures: fidelity+, fidelity-, and conciseness. Each subgraph is high-quality, diverse, and non-dominated in at least one measure that is higher than others in the set. This result provides a more comprehensive and intuitive interpretation to explain “why” $v_{t}$ is identified as “illicit” by $\mathsf{GNN}$-based classification. Indeed, $G_{\zeta}^{5}$, $G_{\zeta}^{6}$, and $G_{\zeta}^{7}$ capture different money laundering scenarios: Peel Chain [16], Spindle [15], and a combination, respectively. Therefore, such a skyline explanation supports law agencies by revealing a more comprehensive interpretation of the decision-making of $\mathsf{GNNs}$ on $v_{t}$. $\Box$
@@
We advocate for developing a $\mathsf{GNN}$ explainer that can efficiently generate skyline explanations for large-scale $\mathsf{GNN}$-based analysis. Such an explainer should: (1) generate skyline explanations for designated output of interest and any user-defined set of explanatory measures; and (2) generate a diversified set of skyline explanations upon request; and (3) ensure desirable guarantees in terms of Pareto-optimality. The need for skyline explanations are evident for trustworthy and multifaceted analysis and decision making, as observed in, e.g., drug repurposing [25], cybersecurity analysis [26], fraud detection [27], social recommendation [28], among others, where $\mathsf{GNN}$ output should be clarified from multiple, comprehensive aspects, rather than one-sided, biased perspectives.
@@
Graph Neural Networks. GNNs have demonstrated themselves as powerful tools in performing various graph learning tasks. Recent studies proposed multiple variants of the GNNs, such as graph convolution networks (GCNs) [29], graph attention networks (GATs) [30], Graph Isomorphism Networks (GINs) [31]. These methods generally follow an information aggregation scheme where features of a node are obtained by aggregating features from its neighboring nodes.
@@
Explanation of GNNs. Several $\mathsf{GNN}$ explanation approaches have been studied. (1) Learning-based methods aim to learn substructures of underlying graphs that contribute to the output of a $\mathsf{GNN}$. GNNExplainer [6] identifies subgraphs with node features that maximize the influence on the prediction by learning continuous soft masks for both adjacency matrix and feature matrix. CF-GNNExplainer [7] learns the counterfactual subgraphs, which lead to significant changes of the output if removed from the graphs. PGExplainer [10] parameterizes the learning process of mask matrix using a multi-layer perceptron. GraphMask [32] learns to mask the edges through each layer of $\mathsf{GNNs}$ that leads to the most sensitive changes to their output. (2) Learning-based $\mathsf{GNN}$ explainers require prior knowledge of model parameters and incur learning overhead for large graphs. Post-hoc approaches perform post-processing to directly compute subgraphs that optimize a fixed criteria. For example,  [12] utilizes Monte-Carlo tree search to compute subgraphs that optimize a game-theory-inspired Shapley value.  [33] follows a similar approach, yet optimizes HN values, a topology-aware variant of Shapley value. SAME [34] extends structure-aware Shapley explanations by using expansion-based Monte Carlo tree search to extract multi-grained connected substructures. These methods are constrained to pre-defined criterion, and cannot easily adapt to customizable, multiple explanability criteria. Beyond post-hoc explainers, GSAT [35] uses stochastic attention within the information bottleneck framework to filter irrelevant components, while PGIB [36] combines prototype learning with the bottleneck to highlight prototypical subgraphs.
@@
Closer to our work are $\mathsf{GNN}$ explainers that optimize pre-defined, multiple criteria. CF2 [8] optimizes a linear function of factual and counterfactual measures and learns feature and edge masks as explanations. RoboGExp [37] generates factual, counterfactual, and robust subgraphs, which remain unchanged upon a bounded number of edge modifications. GEAR [9] learns $\mathsf{GNN}$ explainers by adjusting the gradients of multiple objectives geometrically during optimization. $\mathsf{MOExp}$  [11] solves a bi-objective optimization problem to find Pareto optimal set over “simulatability” (factual) and “counterfactual relevance”. Despite these methods generating explanations over multiple criteria, the overall goal remains pre-defined – and not configurable as needed. In addition, $\mathsf{MOExp}$ does not control the size of explanations, which may result in large number of explanations that are hard to inspect.
@@ II Graphs and GNN Explanation
Graph Neural Networks. $\mathsf{GNNs}$ [29, 38, 39] comprise a well-established family of deep learning models tailored for analyzing graph-structured data. $\mathsf{GNNs}$ generally employ a multi-layer message-passing scheme as shown in Equation 1.
@@
Node Classification. Node classification is a fundamental task in graph analysis [40]. A $\mathsf{GNN}$-based node classifier learns a $\mathsf{GNN}$ ${\mathcal{M}}:(X,\mathbf{A})\rightarrow\mathbf{Y}$ s.t. ${\mathcal{M}}(v)=y_{v}$ for $v\in V_{Tr}\subseteq V$, where $V_{Tr}$ is the training set of nodes with known (true) labels $Y_{Tr}$. The inference process of a trained $\mathsf{GNN}$ $\mathcal{M}$} assigns the label for a test node $v_{t}\subseteq V_{T}$, which are derived from their computed embeddings.
@@
A $\mathsf{GNN}$ explainer is an algorithm that generates explanatory subgraphs in $\zeta$ for ${\mathcal{M}}(v,G)$. An explainability measure $\phi$ is a function: $G_{\zeta}\rightarrow$$\mathbf{R}$ that associates an explanatory subgraph with an explainability score. Given $G$ and ${\mathcal{M}}(v,G)$ to be explained, existing explainers typically solve a single-objective optimization problem: $G_{\zeta}^{*}$ = $\operatorname*{arg\,max}_{G_{\zeta}\in\zeta}$ $\phi(G_{\zeta})$.
@@
Skyline Query. A skyline query is a multi-criteria database query that retrieves a set of non-dominated records from a dataset, based on a given set of attributes [41, 42]. A record dominates another if it is at least as good in all dimensions and strictly better in at least one. For example, when searching for hotels, a skyline query would return those that are not worse than any other option in both price and distance to the city center, while being strictly better in at least one of them. The result forms the Pareto frontier of the dataset, representing the most interesting or optimal choices without requiring users to specify explicit weighting functions across multiple criteria. Skyline queries are widely used in decision making, recommender systems, and multi-objective optimization tasks.

### v4

@@ I Introduction
Graph models such as graph neural networks ($\mathsf{GNNs}$) have been trained and routinely queried to perform network analysis in, e.g., biochemistry, social and financial networks [Wang2023, you2018graph, cho2011friendship, wei2023neural], among other graph analytical tasks. Given a graph $G$ and a set of test nodes $V_{T}$ in $G$, a $\mathsf{GNN}$ $\mathcal{M}$} can be considered as a function that converts a feature representation of $G$ in the form of $(X,A)$, where $X$ (resp. $A$) refers to the node feature matrix (resp. normalized adjacency matrix) of $G$, to an output representation $Z$ (“output embedding matrix”). The output $Z$ can be post-processed to task-specific output such as labels for classification. An “inference query” invokes the above process to query $\mathcal{M}$}, to perform the inference analysis of $G$.
Despite the promising performance of $\mathsf{GNNs}$, it remains desirable yet nontrivial to interpret their outputs to help users understand $\mathsf{GNN}$-based decision making [yuan2022explainability]. Several methods (“explainers”) have been proposed  [ying2019gnnexplainer, lucic2022cf, tan2022learning, zhang2024gear, luo2020parameterized]. Typically, a $\mathsf{GNN}$ explainer extracts a subgraph $G_{\zeta}$ of $G$ that can best clarify the output of an inference query posed on $\mathcal{M}$} over $G$. This is often addressed by discovering a subgraph $G_{\zeta}$ subject to a pre-defined metric, which quantifies the explainability of $G_{\zeta}$ for the output (as summarized in Table I).
For example, a subgraph $G_{\zeta}$ of a graph $G$ is a “factual” explanation for $\mathcal{M}$} over $G$, if it preserves the output of $\mathcal{M}$} (hence is “faithful” to the output of $\mathcal{M}$}[ying2019gnnexplainer, luo2020parameterized, liu2021multi, tan2022learning]. $G_{\zeta}$ is a “counterfactual” explanation, if removing the edges of $G_{\zeta}$ from $G$ leads to a change of the output of $\mathcal{M}$} on the remaining graph (denoted as $G\setminus G_{\zeta}$[lucic2022cf, liu2021multi, tan2022learning]. Other metrics include $\mathsf{fidelity}$$^{-}$ [lucic2022cf, liu2021multi, tan2022learning] (resp. $\mathsf{fidelity}$$^{+}$ [luo2020parameterized, yuan2021explainability]), which quantifies the explainability of $G_{\zeta}$ in terms of the closeness between the task-specific output of $\mathcal{M}$}, such as the probability of label assignments, over $G_{\zeta}$ (resp. $G\setminus G_{\zeta}$) and their original counterpart $G$, and “conciseness (sparsity)” that favors small explanatory subgraphs.
@@
Figure 1: A Bitcoin transaction network with a target IP address (test node $v_{t}$) that has a label “illicit” to be explained [ellipticpp].
@@
$\mathsf{fac}$ factual ${\mathcal{M}}(v,G)$ = ${\mathcal{M}}(v,G_{\zeta})?$ $\mathsf{true}$$\mathsf{false}$ } a Boolean function  [ying2019gnnexplainer, luo2020parameterized, liu2021multi, tan2022learning]
$\mathsf{cfac}$ counterfactual ${\mathcal{M}}(v,G)\neq{\mathcal{M}}(v,G\setminus G_{\zeta})?$ $\mathsf{true}$$\mathsf{false}$ } a Boolean function  [lucic2022cf, liu2021multi, tan2022learning]
$\mathsf{fdl}$$^{+}$ fidelity+ $Pr({\mathcal{M}}(v,G))-Pr({\mathcal{M}}(v,G\setminus G_{\zeta}))$ [1,1][-1,1] the larger, the better  [chen2024view, yuan2021explainability]
$\mathsf{fdl}$$^{-}$ fidelity- $Pr({\mathcal{M}}(v,G))-Pr({\mathcal{M}}(v,G_{\zeta}))$ [1,1][-1,1] the smaller, the better  [chen2024view, luo2020parameterized]
$\mathsf{conc}$ conciseness $\frac{1}{N}\sum_{i=1}^{N}(1-\frac{|E(G_{\zeta})|}{|E(G)|})$ $[0,1]$ the smaller, the better  [luo2020parameterized, yuan2021explainability]
$\mathsf{shapley}$ Shapley value $\phi(G_{\zeta})=\sum_{S\subseteq P\setminus\{G_{\zeta}\}}\frac{|S|!(|P|-|S|-1)!}{|P|!}m(S,G_{\zeta})$ [1,1][-1,1] total contribution of nodes in $G_{\zeta}$  [yuan2021explainability]
@@
A law enforcement agency has posed an inference query over $\mathcal{M}$} to get the output, and wants to further understand “why” the $\mathsf{GNN}$ asserts $v_{t}$ as an illicit account address. They may further ask clarifications by asking “which fraction of the graph $G$ is responsible for the $\mathsf{GNN}$’s decision on assigning the label “illicit” to the account $v_{t}$?”, and ground this output by referring to known real-world money laundering scenarios [spindle, peel]). For example, “Spindle” [spindle] suggests that perpetrators generate multiple shadow addresses to transfer small amounts of assets along lengthy paths to a specific destination; and Peel Chain [peel] launders large amounts of cryptocurrency through sequences of small transactions, where minor portions are ‘peeled’ from the original address and sent for conversion into fiat currency to minimize the risk.
An explanatory subgraph $G_{\zeta}^{1}$ is a factual explanatory subgraph generated by the explainer in [ying2019gnnexplainer]; and $G_{\zeta}^{2}$ is a counterfactual one generated by [lucic2022cf]. Comparing $G_{\zeta}^{1}$ with $G_{\zeta}^{2}$, we observe the following. (1) $G_{\zeta}^{1}$ is a small and concise explanation that only involves $v_{t}$ and most of its neighbors, which includes neighborhood features that preserve the output of a $\mathsf{GNN}$’s output; nevertheless, it misses important nodes that have a higher influence on $\mathsf{GNN}$’s output, that are not necessarily in $v_{t}$’s direct neighborhood. (2) Such nodes can be captured by a counterfactual explanatory subgraph, as depicted in $G_{\zeta}^{2}$. However, a larger fraction of $G$ is included to ensure that the “removal” of edges incurs enough impact to change the output of the $\mathsf{GNN}$, hence sacrificing “conciseness”, conflicting with users’ need on quick generation of small evidence which is “faithful” to the original $\mathsf{GNN}$ output [liu2021multi].
@@
Can we generate explanations that simultaneously address multiple explainability metrics? A quick solution is to compute subgraphs that optimize a weighted linear combination of all metrics [marler2010weighted, das1997closer]. However, this may lead to a marginally optimal answer over all measurements, overlooking other high-quality and diverse solutions, hence an overkill.
@@
Example 2: Consider another two explanatory subgraphs: $G_{\zeta}^{3}$ is the explanation generated by [luo2020parameterized], an explainer that optimizes both conciseness and factual measures [luo2020parameterized]; $G_{\zeta}^{4}$ is from an explainer that linearly combines factual and counterfactual measures into a single, bi-criteria objective function to be optimized [tan2022learning]. Such methods enforce to optimize potentially “conflicting” measures, or highly correlated ones, either may result in lower-quality solutions that are sensitive to data bias. For example, as fidelity- and conciseness both encourage smaller and faithful explanations, explanatory subgraphs obtained by  [luo2020parameterized] such as $G_{\zeta}^{3}$ turns out to be relatively much smaller and less informative, hardly providing sufficient real-world evidence that can be used to interpret money laundering behaviors. On the other hand, explanations from [tan2022learning] such as $G_{\zeta}^{4}$ may easily be a one-sided, factual or counterfactual evidence. We found that such explanations capture only one type of money laundering scenario at best in most cases. $\Box$
@@
Skyline queries [sky_opr, skylinequery3, skylinequery4, lin2006selecting, KungLP75] compute Pareto sets (“skylines”) that dominate the rest data points across a set of quality measures. Skylines generally offer better and more comprehensive solutions against the aforementioned alternatives [das1997closer, sharma2022comprehensive]. We advocate approaching $\mathsf{GNN}$ explanation by generating a set of subgraphs that are Poreto sets over multiple user-defined explanatory measures, referred to as “skyline explanations”.
@@
Example 3: Consider a “skyline query” that seeks subgraphs addressing multiple explainability criteria, which returns a set of subgraphs $G_{\zeta}^{5}$, $G_{\zeta}^{6}$, and $G_{\zeta}^{7}$. These explanatory subgraphs are selected as a Pareto-optimal set across three explanatory measures: fidelity+, fidelity-, and conciseness. Each subgraph is high-quality, diverse, and non-dominated in at least one measure that is higher than others in the set. This result provides a more comprehensive and intuitive interpretation to explain “why” $v_{t}$ is identified as “illicit” by $\mathsf{GNN}$-based classification. Indeed, $G_{\zeta}^{5}$, $G_{\zeta}^{6}$, and $G_{\zeta}^{7}$ capture different money laundering scenarios: Peel Chain [peel], Spindle [spindle], and a combination, respectively. Therefore, such a skyline explanation supports law agencies by revealing a more comprehensive interpretation of the decision-making of $\mathsf{GNNs}$ on $v_{t}$. $\Box$
@@
We advocate for developing a $\mathsf{GNN}$ explainer that can efficiently generate skyline explanations for large-scale $\mathsf{GNN}$-based analysis. Such an explainer should: (1) generate skyline explanations for designated output of interest and any user-defined set of explanatory measures; and (2) generate a diversified set of skyline explanations upon request; and (3) ensure desirable guarantees in terms of Pareto-optimality. The need for skyline explanations are evident for trustworthy and multifaceted analysis and decision making, as observed in, e.g., drug repurposing [pushpakom2019drug], cybersecurity analysis [cyber], fraud detection [aml1], social recommendation [socialRecom], among others, where $\mathsf{GNN}$ output should be clarified from multiple, comprehensive aspects, rather than one-sided, biased perspectives.
@@
Graph Neural Networks. GNNs have demonstrated themselves as powerful tools in performing various graph learning tasks. Recent studies proposed multiple variants of the GNNs, such as graph convolution networks (GCNs) [kipf2016semi], graph attention networks (GATs) [gat], Graph Isomorphism Networks (GINs) [gin]. These methods generally follow an information aggregation scheme where features of a node are obtained by aggregating features from its neighboring nodes.
@@
Explanation of GNNs. Several $\mathsf{GNN}$ explanation approaches have been studied. (1) Learning-based methods aim to learn substructures of underlying graphs that contribute to the output of a $\mathsf{GNN}$. GNNExplainer [ying2019gnnexplainer] identifies subgraphs with node features that maximize the influence on the prediction by learning continuous soft masks for both adjacency matrix and feature matrix. CF-GNNExplainer [lucic2022cf] learns the counterfactual subgraphs, which lead to significant changes of the output if removed from the graphs. PGExplainer [luo2020parameterized] parameterizes the learning process of mask matrix using a multi-layer perceptron. GraphMask [schlichtkrull2020interpreting] learns to mask the edges through each layer of $\mathsf{GNNs}$ that leads to the most sensitive changes to their output. (2) Learning-based $\mathsf{GNN}$ explainers require prior knowledge of model parameters and incur learning overhead for large graphs. Post-hoc approaches perform post-processing to directly compute subgraphs that optimize a fixed criteria. For example,  [yuan2021explainability] utilizes Monte-Carlo tree search to compute subgraphs that optimize a game-theory-inspired Shapley value.  [zhang2022gstarx] follows a similar approach, yet optimizes HN values, a topology-aware variant of Shapley value. SAME [ye2023same] extends structure-aware Shapley explanations by using expansion-based Monte Carlo tree search to extract multi-grained connected substructures. These methods are constrained to pre-defined criterion, and cannot easily adapt to customizable, multiple explanability criteria. Beyond post-hoc explainers, GSAT [miao2022interpretable] uses stochastic attention within the information bottleneck framework to filter irrelevant components, while PGIB [seo2023interpretable] combines prototype learning with the bottleneck to highlight prototypical subgraphs.
@@
Closer to our work are $\mathsf{GNN}$ explainers that optimize pre-defined, multiple criteria. CF2 [tan2022learning] optimizes a linear function of factual and counterfactual measures and learns feature and edge masks as explanations. RoboGExp [robogexp] generates factual, counterfactual, and robust subgraphs, which remain unchanged upon a bounded number of edge modifications. GEAR [zhang2024gear] learns $\mathsf{GNN}$ explainers by adjusting the gradients of multiple objectives geometrically during optimization. $\mathsf{MOExp}$  [liu2021multi] solves a bi-objective optimization problem to find Pareto optimal set over “simulatability” (factual) and “counterfactual relevance”. Despite these methods generating explanations over multiple criteria, the overall goal remains pre-defined – and not configurable as needed. In addition, $\mathsf{MOExp}$ does not control the size of explanations, which may result in large number of explanations that are hard to inspect.
@@ II Graphs and GNN Explanation
Graph Neural Networks. $\mathsf{GNNs}$ [kipf2016semi, du2021multi, jiang2025ICML] comprise a well-established family of deep learning models tailored for analyzing graph-structured data. $\mathsf{GNNs}$ generally employ a multi-layer message-passing scheme as shown in Equation 1.
@@
Node Classification. Node classification is a fundamental task in graph analysis [nc]. A $\mathsf{GNN}$-based node classifier learns a $\mathsf{GNN}$ ${\mathcal{M}}:(X,\mathbf{A})\rightarrow\mathbf{Y}$ s.t. ${\mathcal{M}}(v)=y_{v}$ for $v\in V_{Tr}\subseteq V$, where $V_{Tr}$ is the training set of nodes with known (true) labels $Y_{Tr}$. The inference process of a trained $\mathsf{GNN}$ $\mathcal{M}$} assigns the label for a test node $v_{t}\subseteq V_{T}$, which are derived from their computed embeddings.
@@
A $\mathsf{GNN}$ explainer is an algorithm that generates explanatory subgraphs in $\zeta$ for ${\mathcal{M}}(v,G)$. An explainability measure $\phi$ is a function: $G_{\zeta}\rightarrow{\mathcal{R}}$ that associates an explanatory subgraph with an explainability score. Given $G$ and ${\mathcal{M}}(v,G)$ to be explained, existing explainers typically solve a single-objective optimization problem: $G_{\zeta}^{*}$ = $\operatorname*{arg\,max}_{G_{\zeta}\in\zeta}$ $\phi(G_{\zeta})$.
@@
Skyline Query. A skyline query is a multi-criteria database query that retrieves a set of non-dominated records from a dataset, based on a given set of attributes [papadias2003optimal, papadias2005progressive]. A record dominates another if it is at least as good in all dimensions and strictly better in at least one. For example, when searching for hotels, a skyline query would return those that are not worse than any other option in both price and distance to the city center, while being strictly better in at least one of them. The result forms the Pareto frontier of the dataset, representing the most interesting or optimal choices without requiring users to specify explicit weighting functions across multiple criteria. Skyline queries are widely used in decision making, recommender systems, and multi-objective optimization tasks.